OpenStudy (anonymous):

f(x) = x + 4, find f(3). i did 3X=X+4 -1X = -1X +4 2X = 4 2X/2 =4/2 X = 2 what did i do wrong ?

OpenStudy (unklerhaukus):

f(x) = x + 4, f(3) = (3) + 4 = 7

OpenStudy (anonymous):

so its saying the 3 goes in for the x not f?

OpenStudy (unklerhaukus):

yeah the 3 is playing the role of the x

OpenStudy (anonymous):

ohh so for \[4(x)=x^{2} +3 g(4) \] the answer would be 19 ?